OpenStudy (anonymous):

(1)/(1+(square root of 3)-(square root of 5))

OpenStudy (anonymous):

yeah.. it's a bunch of numbers.. then what? lol

OpenStudy (dumbcow):

\[\frac{1}{1+\sqrt{3}-\sqrt{5}}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i need to use the conjugate to rid of the square roots on the denominator

OpenStudy (dumbcow):

multiply top and bottom by conjugate conjugate is \[1-(\sqrt{3}-\sqrt{5})\]

OpenStudy (anonymous):

well.. the answer is not that easy: \[[(2\sqrt{3}+1)\sqrt{5}+3\sqrt{3}+7]/11\]

OpenStudy (anonymous):

wouldn't it be \[1-\sqrt{3}+\sqrt{5}\]

OpenStudy (dumbcow):

yes same thing, i just didnt distribute the negative

OpenStudy (dumbcow):

however after you multiply them you should get \[-7 +2\sqrt{15}\] so you have to repeat the process and multiply by its conjugate on top and bottom

OpenStudy (anonymous):

can you help by showing the steps, i'm totally getting the wrong answer, i'm making it way too complicated

OpenStudy (curry):

hey dumb cow pleas ehelp i really need your help

OpenStudy (dumbcow):

\[\frac{1}{1+\sqrt{3}-\sqrt{5}}*\frac{1-\sqrt{3}+\sqrt{5}}{1-\sqrt{3}+\sqrt{5}} = \frac{1-\sqrt{3}+\sqrt{5}}{1-\sqrt{3}+\sqrt{5}+\sqrt{3}-3+\sqrt{15}-\sqrt{5}+\sqrt{15}-5} = \frac{1-\sqrt{3}+\sqrt{5}}{-7 +2\sqrt{15}}\]

OpenStudy (dumbcow):

\[=\frac{1-\sqrt{3}+\sqrt{5}}{-7 +2\sqrt{15}}\]

OpenStudy (anonymous):

yessss

OpenStudy (anonymous):

but i have to work with that conjugate right?

OpenStudy (dumbcow):

\[\frac{1-\sqrt{3}+\sqrt{5}}{-7 +2\sqrt{15}}*\frac{-7 -2\sqrt{15}}{-7 -2\sqrt{15}}\]

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

thank you soo much

OpenStudy (anonymous):

can you keep going

OpenStudy (anonymous):

please